OpenStudy (anonymous):

Hey who wants a medal?? Help me show my work with this!! Please?? I need to learn this stuff.

OpenStudy (anonymous):

Suppose a parabola has vertex (–4, 7) and also passes through the point (–3, 8). Write the equation of the parabola in vertex form.

OpenStudy (anonymous):

@UsArmy3947

OpenStudy (anonymous):

@Nutcase482

OpenStudy (henryblah):

Do you know the vertex form in the first place

OpenStudy (anonymous):

Vertex form is: a(x-h)^2+k

OpenStudy (anonymous):

I don't understand this stuff from the very beginning though.

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=V-QVQ7YhQNE

OpenStudy (anonymous):

I cant use youtube.

OpenStudy (henryblah):

The goal is to get the right y value when you put in the x value for the vertex. y=a(x-h)^2+k. Let's say the vertex is (1,1) Then the vertex equation would be y=a(x-1)^2+1 because when you put x=1 into the equation the first part neatly cancels and you are left with y=1. Try and do this with your equation.

OpenStudy (jdoe0001):

hmmm do you even know what a parabola is?

OpenStudy (anonymous):

Yeah..

OpenStudy (anonymous):

There's just too much to learn with this stuff.. Its stupid.

OpenStudy (jdoe0001):

http://www.mathwarehouse.com/geometry/parabola/images/standard-vertex-forms.png well, notice the equation of a parabola in standard form, and in vertex form you'rea sked to use the "vertex form" for this matter how would it look like given the vertex is at ( -4, 7) ?

OpenStudy (anonymous):

Well it would be in a U shape..

OpenStudy (jdoe0001):

hmmm ... I meant how would the equation look like